OpenStudy (anonymous):

Is (x+3)^2 -4=2 is Linear or Quadratic Equation ?

OpenStudy (anonymous):

It has highest power = 2 so what do you think??

OpenStudy (unklerhaukus):

neither the solution is a point

OpenStudy (anonymous):

Quadratic ?

OpenStudy (anonymous):

Then it is...

OpenStudy (unklerhaukus):

ops two points

OpenStudy (anonymous):

You are right @kristine_08

OpenStudy (anonymous):

:)

OpenStudy (unklerhaukus):

\[x=-3\pm\sqrt6\], how can this be a quadratic ?

OpenStudy (anonymous):

She is asking about the equation and not the solution @UnkleRhaukus

OpenStudy (unklerhaukus):

the equations hold the same information

OpenStudy (anonymous):

Your solution is linear but the equation is quadratic having degree = 2.. And you too have found two values of x.. @UnkleRhaukus

OpenStudy (unklerhaukus):

there is only one variable

OpenStudy (anonymous):

\(x=-3 + \sqrt6\) and \(x=-3- \sqrt6\)

OpenStudy (unklerhaukus):

yeah how is that a function ?

OpenStudy (anonymous):

How do you get one value of x??? There are two values of x that you have found..

OpenStudy (unklerhaukus):

i get the same two values as you do

OpenStudy (anonymous):

2(x+3)=4(x+5) Is that a Linear or Quadratic ?

OpenStudy (anonymous):

Yes... Now tell me if an equation has two roots or solution then what the equation called?? Linear Quadratic or cubic???

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I am talking to @UnkleRhaukus

OpenStudy (anonymous):

2(x+3)=4(x+5) Is that a Linear or Quadratic ?

OpenStudy (anonymous):

Firstly tell me what is the highest power of x in this expression?? @kristine_08

OpenStudy (unklerhaukus):